OpenStudy (anonymous):

How can you simplify sqrt(7){sqrt(14) + sqrt(3)}

OpenStudy (anonymous):

\[\sqrt{7}\times \sqrt{14} + \sqrt{3}\] \[\sqrt{98} + \sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{7}\times \sqrt{14}\] Is the same thing as \[\sqrt{7\times14}\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

\[\sqrt{7} ( \sqrt{14} + \sqrt{3}) \] This is how it is set up on my question.

OpenStudy (anonymous):

@Cha1234

OpenStudy (anonymous):

Oh I didn't realize that sorry, well the rule still applies so it would be \[\sqrt{7\times14} + \sqrt{7\times3}\] \[\sqrt{98} + \sqrt{21}\]